OpenStudy (anonymous):

simplify 6x+9/15x^2 divided by 16 x-12/10x^4?

OpenStudy (saifoo.khan):

\[\frac{3(2x+3)}{15x^2} \times \frac{10x^4}{4(4x-3)}\]Can u solve now?